OpenStudy (anonymous):

How to differentiate 2e^y

OpenStudy (anonymous):

\[\frac{ d }{ dy }2e^y = 2e^y\]

OpenStudy (anonymous):

Remember the derivative of e^x = e^x

OpenStudy (anonymous):

Well there's a proof, it requires limits, and \[\frac{ d }{ dx }lnx=\frac{ 1 }{ x }\]

OpenStudy (anonymous):

Basically \[\frac{ d }{ dx } e^x = e^x \implies \frac{ d }{ dx } lnx =\frac{ 1 }{ x }\] \[\frac{ d }{ dx }(\ln(e^x)) = \frac{ d }{ dx }xlne=\frac{ d }{ dx }x =1 \]

OpenStudy (anonymous):

But you can use the chain rule, with the same derivative.

OpenStudy (anonymous):

Simpler, I'd say...

OpenStudy (luigi0210):

\(\Large \color{green}{ლ(́◉◞౪◟◉‵ლ)}\)

OpenStudy (unklerhaukus):

\[f(x)=2e^y\\ f'(x)=2e^y\frac{\mathrm d y}{\mathrm dx}\]